OpenStudy (anonymous):

The equations of two lines are x - 3y = 6 and y = 3x + 2. Determine if the lines are parallel, perpendicular or neither. are the parallel?

OpenStudy (anonymous):

find the slope of each. second has slope 3 because \[y=3x+2\] looks like \[y=mx+b\]

OpenStudy (anonymous):

first has slope \[-3y=-x+6\] \[3y=x-6\] \[y=\frac{1}{3}x-2\]

OpenStudy (aroub):

3y=x+6 , y=1/3x+2 and y=3x+2 so the lines are not parallel nor perpendicular so ill say neither

OpenStudy (anonymous):

they are not parallel because the slopes are not the same, and they are not perpendicular because their product is not -1

OpenStudy (anonymous):

the product of the slopes is 1, so i guess they were trying to trick you into saying perpendicular but they are not